OpenStudy (anonymous):

is (4x)^2 equal to 16x^2?

OpenStudy (anonymous):

U got it

OpenStudy (anonymous):

yes

OpenStudy (nowhereman):

indeed \[(a\cdot b)^c = a^c \cdot b^c\]